OpenStudy (anonymous):

help! Please! glee is on in forty five minutes and i need to finish! 35-5(2x+3)=20 i keep ending up with -10x=0 but that can't be right..... help!

OpenStudy (anonymous):

let's distribute 5 10x+15=20 10x=5 x=5/10 x=1/2

OpenStudy (anonymous):

but its negative five and theres a 35 before it. like 35+-5(2x+3)=20 35+-10x+-15=20 20+-10x=20 -20 -10x=0 then what? what is 0 divided by -10?

OpenStudy (saifoo.khan):

what if you miss today's glee? :O

OpenStudy (anonymous):

HEY! :( its the kick off. comeon!!!!!

OpenStudy (saifoo.khan):

lol. i was jk.

OpenStudy (saifoo.khan):

and the answer is x=0

OpenStudy (anonymous):

then....

OpenStudy (anonymous):

how ?

OpenStudy (saifoo.khan):

-10x=0 x = 0/10 x = 0

OpenStudy (anonymous):

oh okay, thank you so m uch!!!!

OpenStudy (saifoo.khan):

no probs! :)